Write a quadratic quation having the given numbers as solutions.
-5 and -1
The quadratic equation is ? = 0
(Type your answer in standard form. Type an expression using x as a variable.)
Thank very much for your help with this problem.

Answer by ankor@dixie-net.com(22740)

Write a quadratic equation having the given numbers as solutions.
-5 and -1
These solutions would be derived from factors;
(x+5)(x+1) = 0
FOIL
x^2 + x + 5x + 5 = 0
:
x^2 + 6x + 5 = 0